# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v3

@@ Abstract
On tabular data, a significant body of literature has shown that current deep learning (DL) models perform at best similarly to Gradient Boosted Decision Trees (GBDTs), while significantly underperforming them on outlier data (Gorishniy et al., 2021; Rubachev et al., 2022; McElfresh et al., 2023). We identify a natural tabular data setting where DL models can outperform GBDTs: tabular Learning-to-Rank (LTR) under label scarcity. Tabular LTR applications, including search and recommendation, often have an abundance of unlabeled data, and scarce labeled data. We show that DL rankers can utilize unsupervised pretraining to exploit this unlabeled data. In extensive experiments over both public and proprietary datasets, we show that pretrained DL rankers consistently outperform GBDT rankers on ranking metrics—sometimes by as much as $38\%$—both overall and on outliers.
Machine Learning, ICML
@@ 1 Introduction
The learning-to-rank (LTR) problem aims to train a model to rank a set of items according to their relevance or user preference (Liu, 2009). An LTR model is typically trained on a dataset of queries and associated query groups (i.e., a set of potentially relevant documents or items per query), as well as an associated (generally incomplete) ground truth ranking of the items in the query group. The model is trained to output the optimal ranking of documents or items in a query group, given a query. LTR is a core ML technology in many real world applications—most notably in search contexts including Bing web search (Qin & Liu, 2013), Amazon product search (Yang et al., 2022), and Netflix movie recommendations (Lamkhede & Kofler, 2021). More recently, LTR has also found use in Retrieval Augmented Generation (RAG), where a retrieval system is paired with an LLM (large language model) to ground latter’s responses to truth and reduce hallucination (Lewis et al., 2020; Glass et al., 2022; Pan et al., 2022).
@@
Figure 1: Main result (Simulated crowdsourcing of labels): We compare NDCG (normal-$\uparrow$) as we sweep the percentage of labeled training query groups (QGs); labeled QGs have relevance scores for every element of the QG. For small enough fractions of labeled query groups, pre-trained DL rankers (sometimes significantly) outperform both GBDTS and non-pretrained DL rankers. To the left of the black dotted line, pretrained rankers perform the best (statistically significant at $p=0.05$ level). Points are averages over 3 trials.
@@
A natural question is whether deep models can similarly use unsupervised pretraining to outperform GBDTs on the tabular LTR problem. To this end, many papers have studied unsupervised pretraining techniques—both using deep learning and GBDTs—for tabular data (Yoon et al., 2020; Ucar et al., 2021; Bahri et al., 2021; Verma et al., 2021; Majmundar et al., 2022; Duh & Kirchhoff, 2008; Pseudo-Label, 2013). However, none of these methods have shown convincingly that deep learning (with or without unsupervised pretraining) consistently outperforms GBDTs (with or without unsupervised pretraining); at best, deep methods appear to achieve roughly the same performance as GBDTs (Jeffares et al., 2023; Rubachev et al., 2022; Gorishniy et al., 2021; Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022). Furthermore, a recent work showed that GBDTs are significantly more robust against outliers in the dataset than deep learning models (McElfresh et al., 2023).
@@
In this work, we show through extensive empirical evaluation that deep methods with unsupervised pretraining consistently outperform GBDTs—sometimes by a large margin—when labels are scarce. By scarce, we mean that only a small number of query groups contain labels. This setting is standard in most real-world deployments of LTR; for instance, scarcity can arise when organizations: (a) crowdsource labels for a limited number of query groups due to cost constraints, and/or (b) use rare user feedback (e.g., clicks, purchases) as implicit labels. Our experiments consider both sources of scarcity. They are run on the three standard public datasets in the ranking literature: MSLRWEB30K (Qin & Liu, 2013), Yahoo (Chapelle & Chang, 2011), Istella_S (Lucchese et al., 2016), as well as an industry-scale proprietary ranking dataset.
@@
Contributions: (1) We demonstrate for the first time that unsupervised pretraining can produce deep models that outperform GBDTs in ranking. Specifically, pretrained deep rankers can achieve NDCG scores (normalized discounted cumulative gain) (Burges, 2010) up to 38% higher than GBDTs when labels are scarce (Figure 1). For instance, Figure 1 shows on three datasets that as the labels get scarcer, GBDTs start to substantially underperform pretrained deep rankers (experimental setup in Section 4.1.2). Due to the prevalence of scarce labels in practical settings, our results suggest a possible change in how to approach LTR.
(2) We demonstrate that when labels are scarce, pretrained deep rankers can perform significantly better than GBDTs on outlier query groups, boosting NDCG by as much as 80% (Figure 5, Section A.3.4). This is perhaps surprising, given prior results in the label-rich regime showing that GBDTs are significantly more robust to outliers than deep rankers (McElfresh et al., 2023).
@@
(3) We provide empirically-justified LTR-specific pretraining strategies, including a new ranking-specific pretraining loss, SimCLR-Rank. In experiments, SimCLR-Rank achieves similar or better NDCG performance to most existing pretraining methods at an order-of-magnitude speedup, thus enabling scaling to larger models and datasets.
@@
Related work: We focus on the traditional LTR setting where the features are all numeric (tabular data). In this setting, gradient boosted decision trees (GBDTs) (Friedman, 2001) are the de-facto models, and deep models have yet to outperform them convincingly (Qin et al., 2021; Joachims, 2006; Ai et al., 2019; Bruch et al., 2019; Ai et al., 2018; Pang et al., 2020; McElfresh et al., 2023). Self-supervised learning (SSL) or unsupervised pretraining has improved overall performance and robustness to noise (Hendrycks et al., 2019) in settings where there is a significant source of unlabeled data like text (Devlin et al., 2018; Howard & Ruder, 2018; Kitaev et al., 2018; Liu et al., 2017; Song & Roth, 2015) and images (Chen et al., 2020; Chen & He, 2021; Grill et al., 2020; Ge et al., 2021; Xie et al., 2020).
@@ 2 Learning-To-Rank problem and its metrics
In the LTR problem, samples are query groups (QGs) consisting of a query (e.g. shopping query) and $L$ potentially relevant items (e.g. products) for this query. The $i$-th relevant item here is represented by a feature vector $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$, which captures information about this query-item pair, subsuming any query related features. In labeled data, such as training or testing, we also have an associated scalar relevance label $y_{i}$, which could be a binary, ordinal, or real-valued measurement of the relevance of this item for this query (Qin et al., 2021). The objective is to learn a function that ranks these $L$ items such that the items with highest true relevance are ranked at the top.
@@
$\text{NDCG}(\pi_{f_{\theta}},\{y_{i}\}_{i=1}^{L})=\frac{\text{DCG}(\pi_{f_{\theta}},\{y_{i}\}_{i=1}^{L})}{\text{DCG}(\pi^{*},\{y_{i}\}_{i=1}^{L})}\in[0,1],$
where $\pi_{f_{\theta}}:[L]\to[L]$ ($[L]\triangleq\{1,\dots,L\}$) is a ranking of the $L$ elements induced by the scoring function $f_{\theta}$ on $\{\boldsymbol{x}_{i}\}_{i=1}^{L_{k}}$, while $\pi^{*}$ is the ideal ranking induced by the true relevance labels $\{y_{i}\}_{i=1}^{L}$, and discounted cumulative gain (DCG) is defined as $\text{DCG}(\pi,\{y_{i}\}_{i=1}^{L})=\sum_{i=1}^{L}\frac{2^{y_{i}}-1}{\text{log}_{2}(1+\pi(i))}.$ Typically, a truncated version of NDCG is used that only considers the top-$k$ ranked items, denoted as NDCG@$k$. In the rest of our paper, we will refer to NDCG@5 as NDCG; this will be our evaluation metric.
We consider a natural label-scarce LTR setting with large amounts of unlabeled data and a limited supply of labels for a subset of this data. The labels that do exist can arise from multiple sources, such as crowdsourced workers paid to label QGs, or through implicit user feedback (e.g., clicks or purchases); we model both in our evaluation. We use all the query groups for unsupervised pretraining, but we use only the labeled query groups for supervised training.
@@ 2.1 Outlier-NDCG for outlier performance evaluation
Motivated by a recent observation that GBDTs outperform DL models in datasets with irregularities (McElfresh et al., 2023), we also study performance on outliers. In interactive ML systems like search, performing well on outlier queries is particularly valuable as it empowers users to search for more outlier queries, which in turn allows the modeler to collect more data and improve the model. To this end, we evaluate the outlier performance by computing the average NDCG on outlier queries groups and refer to it as Outlier-NDCG for brevity. In practice, some outlier queries may already be known, and the modeler can define the outlier datasets accordingly, like in our proprietary dataset. For example, since industry data pipelines often have missing data/features one could identify samples with missing features as outliers. When outliers are unknown, detecting them (especially in higher-dimensional data) is challenging. Hence, we create a heuristic algorithm by assuming outliers are rare values that are separated from the bulk of the data. Using it we systematically selected outlier query groups for the public datasets that we use. More details about the outliers are provided in Section A.2.2
@@ 3 Design: Unsupervised pretraining for LTR
Semi-supervised learning for GBDTs: While to the best of our knowledge there are no methods to perform unsupervised pretraining in GBDTs, one can use semi-supervised methods like consistency regularization (Jeong & Shin, 2020; Sajjadi et al., 2016; Miyato et al., 2018; Oliver et al., 2018; Berthelot et al., 2019), pseudo-labeling/self-training (Rizve et al., 2021; Pseudo-Label, 2013; Shi et al., 2018; Yarowsky, 1995; McClosky et al., 2006), and PCA (Duh & Kirchhoff, 2008) to address limited size of labeled data. In consistency regularization, one increases the size of the training set by using data augmentations. In pseudo-labeling/self-training, a model trained on the available labeled data is used to label unlabeled data. Then, the final model is trained on the resulting “labeled” dataset. In PCA, one projects the features along directions of maximum variability in the union of labeled and unlabeled datasets. Since a prior work has found that consistency regularization decreases the performance of GBDTs (Rubachev et al., 2022), we evaluate only pseudo-labeling and PCA.
@@
SimCLR (Chen et al., 2020): In a batch of $B$ query-groups, the feature vector $\boldsymbol{x}_{q,i}$ of $q$-th query group and its $i$-th item is stochastically augmented twice to obtain $\boldsymbol{x}_{q,i}^{(0)}$ and $\boldsymbol{x}_{q,i}^{(1)}$, which are called a positive pair. In this paper, we use combinations of the following two augmentation strategies for all domain-agnostic pretraining methods: (a) Zeroing: Zero out features selected randomly and i.i.d. with fixed probability. (b) Gaussian noise: Add Gaussian noise of a fixed scale to all features. Further details are given in Section A.2.3.
@@
Next, a base encoder $h(\cdot)$ and projection head $g(\cdot)$ map $\boldsymbol{x}_{q,i}^{(a)}$ to $\boldsymbol{z}_{q,i}^{(a)}=g(h(\boldsymbol{x}_{q,i}^{(a)}))$ for $a=0,1$. Then by minimizing the InfoNCE loss (Oord et al., 2018) in this projected representation space, we push the positive pair, $(\boldsymbol{z}_{q,i}^{(0)},\boldsymbol{z}_{q,i}^{(1)})$, closer while pushing them farther from all other augmented data points in the batch, which are called negatives. Formally, for $\boldsymbol{x}_{q,i}$, we minimize $\ell_{q,i}^{(0)}+\ell_{q,i}^{(1)}$, where:
$\ell_{q,i}^{(a)}=-\text{cos}(\boldsymbol{z}_{q,i}^{(a)},\boldsymbol{z}_{q,i}^{(\bar{a})})/\tau+\text{log}\sum_{q^{\prime}=1}^{B}\sum_{i^{\prime}=1}^{L_{q^{\prime}}}\sum_{a^{\prime}=0}^{1}\big{[}$ (1)
$\mathbf{1}\{(q^{\prime},i^{\prime},a^{\prime})\neq(q,i,a)\}\cdot\exp(\text{cos}(\boldsymbol{z}_{q,i}^{(a)},\boldsymbol{z}_{q^{\prime},i^{\prime}}^{(a^{\prime})})/\tau)\big{]}\,,$
@@
$\bar{a}=1-a$, $\tau$ is the temperature parameter, $L_{q^{\prime}}$ is the number of items in $q^{\prime}$-th query group, and $\text{cos}(\boldsymbol{z},\widetilde{\boldsymbol{z}})=\langle\boldsymbol{z},\widetilde{\boldsymbol{z}}\rangle/\|\boldsymbol{z}\|\|\widetilde{\boldsymbol{z}}\|$ denotes the cosine similarity. After pretraining, the encoder $h$ is used in downstream applications. SimCLR achieves superior performance in many domains (Chen & He, 2021; Wang et al., 2022b, a). It is known that contrasting a pair of data samples which are hard to distinguish from each other, called hard negatives can help an encoder learn good representations (Robinson et al., 2020; Oh Song et al., 2016; Schroff et al., 2015; Harwood et al., 2017; Wu et al., 2017; Ge, 2018; Suh et al., 2019). SimCLR simply contrasts all pairs of data samples in the batch (including from other query groups) with the assumption that a large enough batch is likely to contain a hard negative.
@@
SimSiam (Chen & He, 2021): Similar to SimCLR, for each data point $\boldsymbol{x}_{q,i}$, this method produces stochastically-augmented positive pairs and their projected representations. However, we pass these representations further through a predictor $\text{pred}(\cdot)$, to get $\boldsymbol{p}_{q,i}^{(a)}=\text{pred}(\boldsymbol{z}_{q,i}^{(a)})$ for $a=0,1$. Finally, we maximize the similarity between the projected and predicted representations: $\text{cos}(\boldsymbol{p}_{q,i}^{(0)},\text{sg}(\boldsymbol{z}_{q,i}^{(1)}))+\text{cos}(\boldsymbol{p}_{q,i}^{(1)},\text{sg}(\boldsymbol{z}_{q,i}^{(0)}))$, where sg is a stop-gradient. Unlike SimCLR, there are no negatives, i.e., the loss function does not push the representation of an augmented sample away from that of other samples’ augmentations. The asymmetry in the loss due to the stop gradient and the predictor prevents representation collapse to make negatives unnecessary (Tian et al., 2021; Zhang et al., 2021; Wang et al., 2022b). If the number of items in each query group is $L$, then the time/space complexity for SimSiam is only $O(BL)$ per batch, whereas it is $O(B^{2}L^{2})$ for SimCLR. Therefore, SimSiam is more efficient and it can scale to larger batchsizes and data.
Figure 2: In SimCLR, each positive pair is contrasted with all other items in the batch. In SimCLR-Rank, each positive pair is contrasted with only the items in the same QG.
Table 1: Pretraining methods: A comparison across unsupervised pretraining methods on NDCG, averaged over 3 trials. Here only 0.2% of the query groups have labels for finetuning. Bold numbers denote the best in a column, and underlined numbers are second best. We find that SimCLR-Rank and SimSiam are the best pretraining methods for tabular LTR. Note that the GBDT methods are deterministic, with standard error 0.
Method MSLR ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
GBDT 0.3908 $\pm$ 0.0000 0.5932 $\pm$ 0.0000 0.5807 $\pm$ 0.0000
GBDT + PCA (Duh & Kirchhoff, 2008) 0.3831 $\pm$ 0.0000 0.5932 $\pm$ 0.0000 0.5758 $\pm$ 0.0000
GBDT + pseudolabeling 0.3773 $\pm$ 0.0000 0.5967 $\pm$ 0.0000 0.5396 $\pm$ 0.0000
SCARF (Bahri et al., 2021) 0.3905 $\pm$ 0.0015 0.5976 $\pm$ 0.0028 0.5792 $\pm$ 0.0037
DACL+ (Verma et al., 2021) 0.3885 $\pm$ 0.0023 0.5770 $\pm$ 0.0049 0.5800 $\pm$ 0.0040
VIME-self (Yoon et al., 2020) 0.3757 $\pm$ 0.0127 0.5386 $\pm$ 0.0229 0.5361 $\pm$ 0.0080
SubTab (Ucar et al., 2021) 0.2889 $\pm$ 0.0025 0.5911 $\pm$ 0.0020 0.4694 $\pm$ 0.0064
SimCLR (Chen et al., 2020) 0.3931 $\pm$ 0.0021 0.6015 $\pm$ 0.0023 0.5828 $\pm$ 0.0047
SimCLR-Rank 0.3959 $\pm$ 0.0022 0.6022 $\pm$ 0.0013 0.5839 $\pm$ 0.0013
SimSiam (Chen & He, 2021) 0.3935 $\pm$ 0.0034 0.6107 $\pm$ 0.0035 0.5618 $\pm$ 0.0049

### v4

@@ Abstract
On tabular data, a significant body of literature has shown that current deep learning (DL) models perform at best similarly to Gradient Boosted Decision Trees (GBDTs), while significantly underperforming them on outlier data Gorishniy et al. (2021); Rubachev et al. (2022); McElfresh et al. (2023). However, these works often study idealized problem settings which may fail to capture complexities of real-world scenarios. We identify a natural tabular data setting where DL models can outperform GBDTs: tabular Learning-to-Rank (LTR) under label scarcity. Tabular LTR applications, including search and recommendation, often have an abundance of unlabeled data, and scarce labeled data. We show that DL rankers can utilize unsupervised pretraining to exploit this unlabeled data. In extensive experiments over both public and proprietary datasets, we show that pretrained DL rankers consistently outperform GBDT rankers on ranking metrics—sometimes by as much as $38\%$—both overall and on outliers.
@@ 1 Introduction
The learning-to-rank (LTR) problem aims to train a model to rank a set of items according to their relevance or user preference (Liu, 2009). An LTR model is typically trained on a dataset of queries and associated query groups (i.e., a set of potentially relevant documents or items per query), as well as an associated (generally incomplete) ground truth ranking of the items in the query group. The model is trained to output the optimal ranking of documents or items in a query group, given a query. LTR is a core ML technology in many real world applications—most notably in search contexts including Bing web search (Qin & Liu, 2013), Amazon product search (Yang et al., 2022), and Netflix movie recommendations (Lamkhede & Kofler, 2021). More recently, LTR has also found use in Retrieval Augmented Generation (RAG), where a retrieval system is paired with an LLM (large language model) to ground the latter’s responses to truth and reduce hallucination Lewis et al. (2020); Glass et al. (2022); Pan et al. (2022).
@@
Figure 1: Main result (Simulated crowdsourcing of labels): We compare NDCG ($\uparrow$) as we sweep the percentage of labeled training query groups (QGs); labeled QGs have relevance scores for every element of the QG. For small enough fractions of labeled query groups, pre-trained DL rankers (sometimes significantly) outperform both GBDTs (represents the best of supervised GBDTs and semi-supervised GBDTs) and non-pretrained DL rankers (represents the best of tabular ResNet (Gorishniy et al., 2021), DeepFM (Guo et al., 2017), and DCNv2 (Wang et al., 2021)). To the left of the black dotted line, pretrained rankers perform the best (statistically significant at $p=0.05$ level). Points are averages over 3 trials.
@@
A natural question is whether deep models can similarly use unsupervised pretraining to outperform GBDTs on the tabular LTR problem. To this end, many papers have studied unsupervised pretraining techniques—both using deep learning and GBDTs—for tabular data Yoon et al. (2020); Ucar et al. (2021); Bahri et al. (2021); Verma et al. (2021); Majmundar et al. (2022); Duh & Kirchhoff (2008); Pseudo-Label (2013). However, none of these methods have shown convincingly that deep learning (with or without unsupervised pretraining) consistently outperforms GBDTs (semi-supervised or supervised); at best, deep methods appear to achieve roughly the same performance as GBDTs (Jeffares et al., 2023; Rubachev et al., 2022; Gorishniy et al., 2021; Shwartz-Ziv & Armon, 2022; Grinsztajn et al., 2022). Furthermore, recent work showed that GBDTs are significantly more robust against outliers in the dataset than deep learning models McElfresh et al. (2023). We note that most of these works study idealized problem settings which may fail to capture complexities of real-world scenarios such as label scarcity and noisy data. We hypothesize that pretrained deep models can beat GBDTs in settings that more accurately reflect real-world challenges.
@@
Through extensive empirical evaluation, we confirm this hypothesis and show that deep models which use unsupervised pretraining consistently outperform GBDTs—sometimes by a large margin—when labels are scarce. By scarce, we mean that only a small number of query groups contain labels. This setting is more standard in most real-world deployments of LTR, rather than the idealized settings studied in prior work. For instance, scarcity can arise when organizations: (a) crowdsource labels for a limited number of query groups due to cost constraints, and/or (b) use rare user feedback (e.g., clicks, purchases) as implicit labels. Our experiments consider both sources of scarcity. They are run on the three standard public datasets in the ranking literature: MSLRWEB30K (Qin & Liu, 2013), Yahoo (Chapelle & Chang, 2011), Istella_S (Lucchese et al., 2016), as well as an industry-scale proprietary ranking dataset.
@@
Contributions: (1) We demonstrate for the first time that unsupervised pretraining can produce deep models that outperform GBDTs in ranking. Specifically, pretrained deep rankers can achieve up to 38% higher NDCG (normalized discounted cumulative gain) score (Burges, 2010) than GBDTs when labels are scarce (Figure 1). For instance, Figure 1 shows on three datasets that as the labels get scarcer, GBDTs start to substantially underperform pretrained deep rankers (experimental setup in Section 4.1.2).
@@
(2) We prescribe empirically-justified LTR-specific pretraining strategies, including a new ranking-specific pretraining loss, SimCLR-Rank, which is a modification of the widely used SimCLR (Chen et al., 2020) loss. In experiments, SimCLR-Rank achieves similar or better performance than most existing pretraining methods at an order-of-magnitude speedup. This allows SimCLR-Rank to scale to large real-world datasets such as our proprietary industry-scale ranking dataset.
(3) We demonstrate that when labels are scarce, pretrained deep rankers can perform significantly better than GBDTs on outlier query groups, boosting NDCG by as much as 80% (Figure 8, Section A.3.4). This is perhaps surprising, given prior results in the label-rich regime showing that GBDTs are significantly more robust to data with irregularities than deep rankers (McElfresh et al., 2023).
The main novelty of this work lies in the identification of label-scarce LTR as a practical real-world tabular learning problem and the results which challenge the prior understanding of the optimal modeling choices for tabular learning. We demonstrate that pretraining strongly help deep models to consistently outperform GBDTs in this label-scarce problem. Of note, these improvements over GBDTs happen even when the labeled data is plentiful in absolute numbers; up to 25% labeled query groups in public datasets and 3 million+ labeled query groups in the large industry-scale ranking dataset. Prior to our work, the ML and ranking literature had not yet identified a widely applicable setting where (pretrained) deep models could consistently outperform GBDTs (Qin et al., 2021; McElfresh et al., 2023). Due to the prevalence of label scarcity in practical settings, our results suggest a positive change in how practitioners may approach real-world tabular learning problems, especially in search and recommendation systems.
@@
Related work: We focus on the traditional LTR setting where the features are all numeric (tabular data). In this setting, gradient boosted decision trees (GBDTs) (Friedman, 2001) are the de-facto models, and deep models have yet to outperform them convincingly (Qin et al., 2021; Joachims, 2006; Ai et al., 2019; Bruch et al., 2019; Ai et al., 2018; Pang et al., 2020; McElfresh et al., 2023). This includes factorization machines (Rendle, 2010), wide & deep architectures (Cheng et al., 2016), and similar architectures designed for recommendation systems (Guo et al., 2017; Wang et al., 2021; Lian et al., 2018; Naumov et al., 2019; Qu et al., 2016). While our paper is architecture agnostic, we perform comparisons of pretrained models against non-pretrained DeepFM and DCNv2 models to show that pretraining is needed to outperform GBDTs consistently. Self-supervised learning (SSL) or unsupervised pretraining has improved overall performance and robustness to noise (Hendrycks et al., 2019) in settings where there is a significant source of unlabeled data like text (Devlin et al., 2018; Howard & Ruder, 2018; Kitaev et al., 2018; Liu et al., 2017; Song & Roth, 2015) and images (Chen et al., 2020; Chen & He, 2021; Grill et al., 2020; Ge et al., 2021; Xie et al., 2020).
@@ 2 Learning-To-Rank problem and its metrics
In the LTR problem, samples are query groups (QGs) consisting of a query (e.g. shopping query) and $L$ potentially relevant items (e.g. products) for this query. The $i$-th relevant item here is represented by a feature vector $\bm{x}_{i}\in\mathbb{R}^{d}$, which captures information about this query-item pair, subsuming any query related features. In labeled data, such as training or testing, we also have an associated scalar relevance label $y_{i}$, which could be a binary, ordinal, or real-valued measurement of the relevance of this item for this query (Qin et al., 2021). The objective is to learn a function that ranks these $L$ items such that the items with highest true relevance are ranked at the top.
@@
$\text{NDCG}(\pi_{f_{\theta}},\{y_{i}\}_{i=1}^{L})=\frac{\text{DCG}(\pi_{f_{\theta}},\{y_{i}\}_{i=1}^{L})}{\text{DCG}(\pi^{*},\{y_{i}\}_{i=1}^{L})}\in[0,1]\,,$ (1)
where $\pi_{f_{\theta}}:[L]\to[L]$ ($[L]\triangleq\{1,\dots,L\}$) is a ranking of the $L$ elements induced by the scoring function $f_{\theta}$ on $\{\bm{x}_{i}\}_{i=1}^{L_{k}}$ (i.e. $\pi(i)$ is the rank of the $i$-th item), while $\pi^{*}$ is the ideal ranking induced by the true relevance labels $\{y_{i}\}_{i=1}^{L}$, and discounted cumulative gain (DCG) is defined as $\text{DCG}(\pi,\{y_{i}\}_{i=1}^{L})=\sum_{i=1}^{L}\frac{2^{y_{i}}-1}{\text{log}_{2}(1+\pi(i))}.$ Typically, a truncated version of NDCG is used that only considers the top-$k$ ranked items, denoted as NDCG@$k$. In the rest of our paper, we will refer to NDCG@5 as NDCG; this will be our evaluation metric.
We consider a natural label-scarce LTR setting with large amounts of unlabeled data and a limited supply of labels for a subset of this data. The labels that do exist can arise from multiple sources, such as crowdsourced workers paid to label QGs, or through implicit user feedback (e.g., clicks or purchases); we model both in our evaluation. Unless specified otherwise, we use all the query groups for unsupervised pretraining, but we use only the labeled query groups for supervised training.
@@ 2.1 Outlier-NDCG for outlier performance evaluation
Motivated by a recent observation that GBDTs outperform DL models in datasets with irregularities McElfresh et al. (2023), we also study their performance on outliers. In interactive ML systems like search, performing well on outlier queries is particularly valuable as it empowers users to search for more outlier queries, which in turn allows the modeler to collect more data and improve the model. To this end, we evaluate the outlier performance by computing the average NDCG on outlier queries groups and refer to it as Outlier-NDCG for brevity. In practice, some outlier queries may already be known, and the modeler can define the outlier datasets accordingly, like in our proprietary dataset. For example, since industry data pipelines often have missing data/features one could identify samples with missing features as outliers. When outliers are unknown, detecting them (especially in higher-dimensional data) is challenging. Hence, we create a heuristic algorithm by assuming outliers are rare values that are separated from the bulk of the data. Using it we systematically selected outlier query groups for the public datasets that we use. More details about the outliers are provided in Section A.2.2. Separately measuring the performance on the outliers is justified by our later observation that performance on the full dataset and the outliers can significantly differ (Tables 2 and 3).
@@ 3 Design: Unsupervised pretraining for LTR
Semi-supervised learning for GBDTs: While to the best of our knowledge there are no methods to perform unsupervised pretraining in GBDTs, one can use semi-supervised methods like consistency regularization (Jeong & Shin, 2020; Sajjadi et al., 2016; Miyato et al., 2018; Oliver et al., 2018; Berthelot et al., 2019), pseudo-labeling/self-training (Rizve et al., 2021; Pseudo-Label, 2013; Shi et al., 2018; Yarowsky, 1995; McClosky et al., 2006), and PCA (Duh & Kirchhoff, 2008) to address limited size of labeled data. In consistency regularization, one increases the size of the training set by using data augmentations. In pseudo-labeling/self-training, a model trained on the available labeled data is used to label unlabeled data. Then, the final model is trained on the resulting “labeled” dataset. In PCA, one projects the features along directions of maximum variability in the union of labeled and unlabeled datasets. Since prior work has found that consistency regularization decreases the performance of GBDTs (Rubachev et al., 2022), we evaluate only pseudo-labeling and PCA.
@@
SimCLR (Chen et al., 2020): In a batch of $B$ query-groups, the feature vector $\bm{x}_{q,i}$ of $q$-th query group and its $i$-th item is stochastically augmented twice to obtain $\bm{x}_{q,i}^{(0)}$ and $\bm{x}_{q,i}^{(1)}$, which are called a positive pair. In this paper, we use combinations of the following two augmentation strategies for all domain-agnostic pretraining methods: (a) Zeroing: Zero out features selected randomly and i.i.d. with fixed probability. (b) Gaussian noise: Add Gaussian noise of a fixed scale to all features. Further details are given in Section A.2.3.
@@
Next, a base encoder $h(\cdot)$ and projection head $g(\cdot)$ map $\bm{x}_{q,i}^{(a)}$ to $\bm{z}_{q,i}^{(a)}=g(h(\bm{x}_{q,i}^{(a)}))$ for $a=0,1$. Then by minimizing the InfoNCE loss (Oord et al., 2018) in this projected representation space, we push the positive pair, $(\bm{z}_{q,i}^{(0)},\bm{z}_{q,i}^{(1)})$, closer while pushing them farther from all other augmented data points in the batch, which are called their negatives. Formally, for $\bm{x}_{q,i}$, we minimize $\ell_{q,i}^{(0)}+\ell_{q,i}^{(1)}$, where:
$\ell_{q,i}^{(a)}=-\text{cos}(\bm{z}_{q,i}^{(a)},\bm{z}_{q,i}^{(\bar{a})})/\tau+\text{log}\sum_{q^{\prime}=1}^{B}\sum_{i^{\prime}=1}^{L_{q^{\prime}}}\sum_{a^{\prime}=0}^{1}\big{[}\mathbf{1}\{(q^{\prime},i^{\prime},a^{\prime})\neq(q,i,a)\}\cdot\exp(\text{cos}(\bm{z}_{q,i}^{(a)},\bm{z}_{q^{\prime},i^{\prime}}^{(a^{\prime})})/\tau)\big{]}$ (2)
@@
$\bar{a}=1-a$, $\tau$ is the temperature parameter, $L_{q^{\prime}}$ is the number of items in $q^{\prime}$-th query group, and $\text{cos}(\bm{z},\widetilde{\bm{z}})=\langle\bm{z},\widetilde{\bm{z}}\rangle/\|\bm{z}\|\|\widetilde{\bm{z}}\|$ denotes the cosine similarity. After pretraining, the encoder $h$ is used in downstream applications. SimCLR achieves superior performance in many domains (Chen & He, 2021; Wang et al., 2022b; a). It is known that contrasting a pair of data samples which are hard to distinguish from each other, called hard negatives can help an encoder learn good representations (Robinson et al., 2020; Oh Song et al., 2016; Schroff et al., 2015; Harwood et al., 2017; Wu et al., 2017; Ge, 2018; Suh et al., 2019). SimCLR simply contrasts all pairs of data samples in the batch (including from other query groups) with the assumption that a large enough batch is likely to contain a hard negative.
@@
SimSiam (Chen & He, 2021): Similar to SimCLR, for each data point $\bm{x}_{q,i}$, this method produces stochastically-augmented positive pairs and their projected representations. However, we pass these representations further through a predictor $\text{pred}(\cdot)$, to get $\bm{p}_{q,i}^{(a)}=\text{pred}(\bm{z}_{q,i}^{(a)})$ for $a=0,1$. Finally, we maximize the similarity between the projected and predicted representations: $\text{cos}(\bm{p}_{q,i}^{(0)},\text{sg}(\bm{z}_{q,i}^{(1)}))+\text{cos}(\bm{p}_{q,i}^{(1)},\text{sg}(\bm{z}_{q,i}^{(0)}))$, where sg is a stop-gradient. Unlike SimCLR, there are no negatives, i.e., the loss function does not push the representation of an augmented sample away from that of other samples’ augmentations. The asymmetry in the loss due to the stop gradient and the predictor prevents representation collapse to make negatives unnecessary Tian et al. (2021); Zhang et al. (2021); Wang et al. (2022b). If the number of items in each query group is $L$, then the time/space complexity for SimSiam is only $O(BL)$ per batch, whereas it is $O(B^{2}L^{2})$ for SimCLR. Therefore, SimSiam is more efficient and it can scale to larger batchsizes and data.
Figure 2: In SimCLR and its variants, positive pairs, or augmented versions of the same data sample, are trained to have similar embeddings. Positive-negative pairs, or augmented samples originating from two different data points or classes, are trained to have distant embeddings. In vanilla SimCLR, each positive pair is contrasted with all other items in the batch, denoted by the data points contained in the brown dashed line above. In SimCLR-Rank, each positive pair is contrasted with only the items in the same Query Group (QG), denoted by the data points inside the red solid line above. By limiting the negatives to a smaller set of harder-to-contrast relevant items, SimCLR-Rank is faster and scales to larger datasets.